OpenStudy (anonymous):

use the natural log to find x... 5^x=2e^x+1... answer in the back of the book is (1+ln2)/(-1+ln5).. but not sure how to get there.. please show work!!!!!!!!

OpenStudy (anonymous):

@phi

OpenStudy (phi):

is the x+1 the exponent ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i had the idea of taking the natural log of both sides

OpenStudy (phi):

that sounds promising. what do you get ?

OpenStudy (anonymous):

ln5^x=2(x+1)

OpenStudy (anonymous):

right?

OpenStudy (phi):

you should do it in steps \[ 5^x=2e^{x+1} \\ \ln\left( 5^x\right) = \ln\left( 2e^{x+1}\right) \]

OpenStudy (anonymous):

cant i just cross out the ln and e on the right though

OpenStudy (phi):

almost. but you have a 2 in there. use log(a*b) = log(a) + log(b) on the right side on the left side use log(a^b) = b*log(a)

OpenStudy (anonymous):

so on left side it would be.. xlog(5)

OpenStudy (phi):

yes. but write out the whole equation.

OpenStudy (anonymous):

im not sure about the right side..

OpenStudy (phi):

use log(a*b) = log(a) + log(b) where a matches 2 and b matches the e^stuff

OpenStudy (anonymous):

okay so...... xlog5=log2+log(e^x+1)

OpenStudy (phi):

and use parens , because (x+1) is the exponent: xlog5=log2+log(e^(x+1) ) now you can "cross out the log and the e"

OpenStudy (anonymous):

ok so.... xlog5=log2+(x+1)

OpenStudy (phi):

yes. solve for x. remember log5 (or maybe we should use ln 5 ? ) is just a constant

OpenStudy (anonymous):

can i subtract the x to the left side of the equation

OpenStudy (phi):

yes

OpenStudy (anonymous):

but wouldnt that cancelout the x's

OpenStudy (phi):

write down what you get

OpenStudy (anonymous):

log5=log2+1

OpenStudy (phi):

xlog5=log2+(x+1) or x log5 = log2 + x + 1 add -x to both sides

OpenStudy (anonymous):

ya and i got log5=log2+1

OpenStudy (phi):

you get x log5 -x = log2 + x + 1 -x on the right side x - x is 0 so x log5 -x = log2 + 1 if this were 1.608*x - x = log2 +1 would the x's disappear ? notice that ln5 is about 1.608...

OpenStudy (phi):

what about this idea? factor an x out of each term on the left side? x ln5 - x = ln2 + 1

OpenStudy (anonymous):

ok then what is is the next step

OpenStudy (anonymous):

divide by ln5?

OpenStudy (phi):

factor an x out of each term on the left side? x ln5 - x = ln2 + 1 what do you get ?

OpenStudy (anonymous):

x(ln5-1)=ln2+1

OpenStudy (phi):

now divide both sides by (ln5 -1)

OpenStudy (anonymous):

sweet thanks so much!